OpenStudy (anonymous):

find the percent of change for: 4Ft to 11 Ft

OpenStudy (bibby):

4*(some percent) = 11 4x=11

OpenStudy (anonymous):

huh

OpenStudy (bibby):

There's a formula for percent increase but that doesn't really matter here. x = 11/4 = 2.75 That's the decimal increase. convert it to percetn

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 4&\to 7\\ &\uparrow \\ &3\ units \end{array}\\ \quad \\ \textit{what's 3 of 4?}\qquad \begin{array}{ccllll} quantity&percentage\\ \hline\\ 4&100\\ 3&x \end{array}\bf \implies \cfrac{4}{3}=\cfrac{100}{x}\)